# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ 5.2 ML-predicted outcome and covariates
Next, we explore a situation where the outcome and covariates are not directly observed in the unlabeled data. Instead, we employ machine learning techniques to predict both. Additionally, we possess a much smaller data with both outcome and covariates measured. This full data can be categorized into two segments: ${\cal L}\cup{\cal U}$. The first segment, ${\cal L}$, includes the data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$, while the second segment, ${\cal U}$, comprises the instances of ${\bf z}_{i}$ for $i=n+1,\cdots,N+n$, and $N>>n$. The variable ${\bf z}$ is to predict both the outcome and the covariates. The ML-predicted outcome and covariates are represented by $\widehat{f}({\bf z})$ and $\widehat{q}({\bf z})$, respectively, where $\widehat{f}$ and $\widehat{q}$ are independently obtained. The POP-Inf estimator for this situation is to solve the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (29)
where $\widehat{\bm{\omega}}$ can be obtained by minimizing the variance of the estimator. Similarly, we have the consistency and asymptotic normality for the $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ in Corollary 4 whose proof follows from the proof of Theorem 1.
Corollary 4.
Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability. Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime\prime}(\bm{\omega})),$ (30)
where ${\bm{\Sigma}}^{\prime\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}^{\prime\prime}_{2}+\rho{\bf M}^{\prime\prime}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}^{\prime\prime}_{4}$, ${\bf M}_{2}^{\prime\prime}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}^{\prime\prime}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.
@@ 6 Simulations
To demonstrate our method and corroborate our theoretical findings, we estimate outcome mean and linear regression coefficients on simulated datasets with the POP-Inf estimator and its competitors. In a variety of simulated datasets, all methods lead to satisfactory coverage. Moreover, the proposed POP-Inf estimator yields the smallest variances. We detail the data generation and implementation in Section 6.1 and present results in Section 6.2.
6.1 Settings
We generate a dataset with $n\in\{250,500\}$ labeled samples and $N=20n$ unlabeled samples, identically and independently drawn from the following data generating process. For unit $i$, we generate ${\bf X}_{i}=[X_{1i},X_{2i}]^{\rm T}$ from a multivariate normal distribution with mean zero and an identity covariance matrix. Without loss of generality, we set $r_{1}=\cdots=r_{n}=1$ and generate their outcome
$Y_{i}=5\gamma\left(X_{1i}+X_{1i}+X_{1i}^{2}+X_{2i}^{2}+X_{1i}X_{2i}\right)/\sqrt{6}+\varepsilon_{i},$ (31)
where $\varepsilon_{i}\sim\mathcal{N}(0,5\sqrt{1-\gamma^{2}})$ independently and $\gamma$ varies between zero and one by a grid of $0.02$. The parameter $\gamma$ controls the data generation and it quantifies the signal to noise ratio. Roughly speaking, the higher $\gamma$ is, the easier it would be to predict $Y$, and henceforth the higher efficiency one would expect for any estimator that utilizes unlabeled data.
For the simplicity of this simulation, we consider ${\bf Z}={\bf X}$. To build the ML prediction algorithm $\widehat{f}$, we generate an independent training set of size $500$ with the same data generating process, on which we train a random forest with R package randomForest (Breiman,, 2001). Arguments have been set to default .
We consider two estimands, the outcome mean and the linear regression coefficient of regressing $Y$ onto $X_{1}$ with independent and identically distributed normal errors. For each estimand, we construct four estimators: the classic estimator, the PP estimator (Angelopoulos et al., 2023a, ), the $*{}^{*}$-based estimator, and the proposed POP-Inf estimator. Simulation results are based on $1000$ repetitions. At each repetition, we calculate standard errors for all estimators based on Theorem 1 and Corollary 2.
@@ 6.2 Results
As shown in the left panels in Figure 1, all methods lead to a coverage rate at around the nominal level $95\%$ when $n=250$, suggesting that they are all unbiased with valid standard error estimates. To compare the efficiency of an estimator compared with the classic approach,, we calculate the ratios of standard error estimates between one estimator and the classic approach. When the ratio is smaller than one, it suggests that the corresponding estimator is more efficient than the classic approach and vice versa. The right panels of Figure 1 plot these ratios against $\gamma\in[0,0.5]$ for POP-Inf, PP, and $*{}^{*}$-based estimators for both estimands at $n=250$. These estimators behave similarly for the two estimands.
When $\gamma$ is close to zero, the POP-Inf estimator is slightly more efficient than the classic approach, while the other two estimators are much less efficient. As $\gamma$ increases to $0.5$, reflecting that $\widehat{f}$ improves, all ratios drop. The POP-Inf estimator always has the lowest ratio, meaning its efficiency improvement against other estimators. When $\gamma\in[0.5,1]$, as can be seen in Figure 2, the POP-Inf estimator still has the smallest standard error, whereas the other two estimators have comparable performance. This suggests that when the prediction performs well, the proposed POP-Inf estimator is still a better choice. Similar patterns can be found for $n=500$ (Figure 4 in Appendix A). In addition, Figure 5 in Appendix A shows boxplots of biases of the POP-Inf estimator across $1000$ repetitions. As the sample size increases, the boxes become narrower and hence the variance becomes smaller. This corroborates the root $n$ consistency.
Figure 2: Relative efficiency in simulations with $\gamma\in[0.5,1]$ and $n=250$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 6. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis. A ratio lower than one represents a smaller standard error than the classic approach.
@@ 7 Data Applications
We used the POP-Inf method to assess the effect of biological sex on gene expression across 44 human tissues using GTEx data described in Example 1. We aim to use a ML algorithm called hypergraph factorization (HYFA) to impute the missing gene expression (Viñas et al.,, 2023) in the uncollected tissue. The imputed gene expression is then used in POP-Inf to increase the statistical power to identify sex-biased genes. We compared POP-Inf with classic, PP, and $*{}^{*}$-based approaches.
@@
In particular, we processed the GTEx-v8 data following the GTEx pipeline (GTEx Consortium et al.,, 2015). We then used HYFA to predict transcript levels for each uncollected tissue for each individual in the data, resulting in a gene expression dataset of $834$ individuals across $44$ tissues. HYFA is a parameter-efficient graph representation learning approach for multi-tissue prediction of gene expression. We used cross-validation to predict transcript levels in the labeled data to avoid over-fitting. Prediction accuracy is measured by the correlation between measured and imputed gene expression in the labeled data. We then used linear regression to assess the effect of sex on gene expression while controlling for technical factors such as surrogate variables, age, RNA integrity number, and total ischemic time for a sample.
Results are shown in Figure 3. Each dot represents the result of the inference for one gene in one tissue in Figure 3 (a)-(f). The point estimates of all method are close to the classic approach (Figure 3 (a)-(c)). However, both PP and $*{}^{*}$-based are less efficient than the classic approach when the prediction accuracy is low. In contrast, the POP-Inf estimator is always no less efficient than the classic approach, regardless of the prediction accuracy (Figure 3 (d)-(f)). Moreover, POP-Inf identifies more sex-biased genes than other approaches, demonstrating its improved efficiency over alternatives (Figure 3 (g)).
In conclusion, the real data results are consistent with our theoretical and simulation results, demonstrating that our method POP-Inf is more attractive in real-world post-prediction inference.
@@
Figure 3: Comparison of POP-Inf, classic, PP, and $*{}^{*}$-based approaches in identifying sex-biased genes using GTEx data. Each panel illustrates a different aspect of comparison on the y- and x- axes: point estimates between the (a) classic and PP approaches; (b) classic and $*{}^{*}$-based approaches; (c) classic and POP-Inf approaches; estimated standard errors of point estimates between the (d) classic and PP approaches; (e) classic and $*{}^{*}$-based approaches; (f) classic and POP-Inf approaches; (g) number of sex-biased genes identified by each of the four approaches. The dashed line represent $y=x$ in (a)-(c) and $y=1$ in (d)-(f).
@@ 8 Discussion
We have provided a simple yet powerful method, the POP-Inf, to improve the efficiency of using unlabeled data with arbitrary ML predictions. To enable inference, we establish the consistency and asymptotic normality of the proposed estimator and prove its superiority over baseline methods. In extensive simulations and real data application, we demonstrate the superiority of the POP-Inf estimator over alternatives.
After finishing this manuscript, we became aware of an independent work by Angelopoulos et al., 2023b that was preprinted in the same month as ours. This work is similar to POP-Inf in using weights to guarantee an efficiency improvement with arbitrary ML predictions. Our development of POP-Inf was conducted independently and without prior knowledge of their work. Moreover, our approach differs in several key aspects. First, we incorporate a weight vector in the the estimating equation, while Angelopoulos et al., 2023b add a scalar weight to the loss function. Therefore, our approach optimizes the weight in a larger set of estimators. Second, we targets an element-wise variance reduction of the (vector) parameter of interest, while Angelopoulos et al., 2023b improve efficiency by reducing the trace of the covariance matrix. Therefore, our method has guaranteed efficiency gain over the classic method for each parameter. This is important in practice, because many statistical analyses have a specific subvector of the parameter garnering scientific attention, whereas other parameters are considered as nuisances (GTEx Consortium,, 2020; Uffelmann et al.,, 2021). Minimizing the trace of the covariance matrix for all elements including those for the nuisance parameters does not guarantee element-wise variance reduction, which can lead to inflated variances for the parameters of interest compared to classic methods. Third, we connect our estimator to the efficient influence function, which is another powerful tool in the pursuit of efficiency. We have demonstrated in numeric comparisons that the POP-Inf estimator is more practical than an $*{}^{*}$-based estimator when a good estimate of the nuisance function cannot be guaranteed.
This work fits into a variety of scientific problems and unlocks several directions of future research. One can apply the same principal to more complicated data structure and missing mechanisms, for example in cases where a multi-dimensional outcome or covariate is subject to missing. It would be an interesting future work to carefully discuss relationships between ${\bf X}$ and ${\bf Z}$ and and how they may affect the magnitude of the efficiency improvement.
The R codes to implement POP-Inf, along with the benchmark methods, and to replicate the simulation and real data analysis, are available at https://github.com/qlu-lab/POPInf.
@@ Acknowledgments
Appendix A Supplementary Figures to Simulations
Figure 4: Relative efficiency and coverage rates in simulations with $\gamma\in[0,1]$ and $n=500$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 6. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis. A ratio lower than one represents a smaller standard error than the classic approach.
Figure 5: Bias of the POP-Inf estimator in simulations for (red) $n=250$ and (purple) $n=500$. Panels are faceted by the estimand and coordinated by $\gamma\in[0,1]$ on the x-axis.

### v4

@@ 5.2 ML-predicted outcome and covariates
Next, we explore a situation where the outcome and covariates are not directly observed in the unlabeled data. Instead, we employ machine learning techniques to predict both. Additionally, we possess much smaller data with both outcome and covariates measured. This full data can be categorized into two segments: ${\cal L}\cup{\cal U}$. The first segment, ${\cal L}$, includes the data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$, while the second segment, ${\cal U}$, comprises the instances of ${\bf z}_{i}$ for $i=n+1,\cdots,N+n$, and $N>>n$. The variable ${\bf z}$ is to predict both the outcome and the covariates. The ML-predicted outcome and covariates are represented by $\widehat{f}$ and $\widehat{q}({\bf z})$, respectively, where $\widehat{f}$ and $\widehat{q}$ are independently obtained. The PSPA estimator for this situation is to solve the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+{\rm diag}({\bm{\omega}})\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f},\widehat{q};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f},\widehat{q};{\bm{\theta}})\right\}={\bf 0},$ (31)
Replacing the estimating equation in Algorithm 2 with Equation 31 gives an algorithm for estimation and statistical inference for this task. The consistency and asymptotic normality of the resulting estimator can be established by the proof of Theorem 1.
@@ 6 Simulations
We conduct simulations to evaluate the performance of PSPA using linear and logistic regression across two settings: one with ML-predicted labels and another with ML-predicted covariates. In the ML-predicted labels setting, we considered classical, PPI, and EIF*-based methods. The EIF*-based method employs a weighting vector defined as ${\bm{\omega}}^{\textnormal{EIF*}}=\left[\frac{N}{N+n},\dots,\frac{N}{N+n}\right]^{\rm T}$, where $N$ and $n$ represent the sample sizes of unlabeled and labeled data, respectively. For the ML-predicted covariates setting, we used classical and imputation-based methods as baselines. We present the implementation details in Section 6.1 and the simulation results in Section 6.2.
6.1 Implementation details
In all simulations, the ground truth coefficients are obtained using $5\times 10^{4}$ samples. The labeled data is with 500 samples and the unlabeled data is with 500, 1500, 2500, 5000, or 10000 samples for different settings. A pre-trained random forest with 100 trees to grow is obtained from a hold-out data with 1000 samples. All simulations are repeated 1000 times.
6.1.1 ML-predicted labels
[…]
6.1.2 ML-predicted covariates
[…]
Figure 2: Coverage of the confidence interval and relative ratio of its width compared to the classical method for linear and logistic regression. ML is used to predict the labels. Panels (a)-(d) show the coverage of the confidence interval. Panels (e)-(h) show the relative ratio of the width of the confidence interval in comparison with the classical method. Panels (a), (b), (e), and (f) correspond to settings with varying sample sizes of unlabeled data. Panels (c), (d), (g), and (h) correspond to settings with different levels of imputation accuracy. The dashed line represent $y=0.95$ in (a)-(d) and $y=1$ in (e)-(h).
@@ 6.2 Results
6.2.1 ML-predicted labels
Figure 2 shows the simulation results for the setting of ML-predicted labels. All evaluated methods achieved confidence interval coverage rates close to 95%, indicating inference validity across different unlabeled sample sizes and different levels of ML prediction accuracy. PPI and EIF*-based methods may be less efficient than classical methods when the ML has poor prediction accuracy or limited unlabeled data. PSPA has narrower confidence intervals compared to classical method and other baseline methods, especially as the unlabeled sample size and ML prediction performance increase.
6.2.2 ML-predicted covariates
Figure 3 shows the simulation results for the setting with ML-predicted covariates. The imputation-based method fails to achieve the correct confidence interval coverage. Both classical and PSPA methods lead to coverage rates around 95%, which suggests valid inference results. PSPA has narrower confidence intervals compared to the classical method, regardless of the unlabeled sample size and the accuracy of the ML model.
Figure 3: Coverage of the confidence interval and relative ratio of its width compared to the classical method for linear and logistic regression. ML is used to predict the covariates. Panels (a)-(d) show the coverage of the confidence interval. Panels (e)-(h) show the relative ratio of the width of the confidence interval in comparison with the classical method. Panels (a), (b), (e), and (f) correspond to settings with varying sample sizes of unlabeled data. Panels (c), (d), (g), and (h) correspond to settings with different levels of imputation accuracy. The dashed line represent $y=0.95$ in (a)-(d) and $y=1$ in (e)-(h).
@@ 7 Data Applications
7.1 Sex-differentiated gene expression
@@
We used the PSPA method to assess the effect of biological sex on gene expression across 44 human tissues using GTEx data. We aim to use an ML algorithm called hypergraph factorization (HYFA) to impute the missing gene expression (Viñas et al.,, 2023) in the uncollected tissue. The imputed gene expression is then used in PSPA to increase the statistical power to identify sex-biased genes. We compared PSPA with classical, PPI, and EIF-based approaches. In particular, we processed the GTEx-v8 data following the GTEx pipeline (GTEx Consortium et al.,, 2015). We then used HYFA to predict transcript levels for each uncollected tissue for each individual in the data, resulting in a gene expression dataset of $834$ individuals across $44$ tissues. HYFA is a parameter-efficient graph representation learning approach for multi-tissue prediction of gene expression. We used cross-validation to predict expression levels in the labeled data to avoid over-fitting. Prediction accuracy is measured by the correlation between measured and imputed gene expression in the labeled data. We then used linear regression to assess the effect of sex on gene expression while controlling for technical factors including surrogate variables, age, RNA integrity number, and total ischemic time for a sample.
@@
Figure 4: Comparison of PSPA, classical, PPI, and EIF-based approaches in identifying sex-biased gene expressions using GTEx data. Each panel illustrates a different aspect of comparison on the y- and x- axes: point estimates between the (a) classical and PPI approaches; (b) classical and EIF-based approaches; (c) classical and PSPA approaches; estimated standard errors between the (d) classical and PPI approaches; (e) classical and EIF-based approaches; (f) classical and PSPA approaches; (g) number of sex-biased genes identified by each of the four approaches. The dashed lines represent $y=x$ in (a)-(c) and $y=1$ in (d)-(f).
Results are shown in Figure 4. Each dot represents the result of the inference for one gene in one tissue in Figure 4 (a)-(f). The point estimates of all methods are close to the classical approach (Figure 4 (a)-(c)). However, both PPI and EIF-based are less efficient than the classical approach when the prediction accuracy is low (Figure 4 (d)-(e)). In contrast, the PSPA estimator is always no less efficient than the classical approach, regardless of the prediction accuracy (Figure 4 (f)). Moreover, PSPA identifies more sex-biased genes than other approaches, demonstrating its improved efficiency over alternatives (Figure 4 (g)).
7.2 Risk factors for bone mineral density
Next, we applied PSPA to a multi-dimensional linear regression task. The goal is to identify the associations between dual-energy x-ray absorptiometry (DXA)-derived total bone mineral density (DXA-BMD) and several covariates, including biological sex, age, physical activities (PA), sedentary behavior (SB), smoking status (current smoker or not), and frequency of alcohol intake. DXA-BMD serves as the primary diagnostic marker for osteoporosis and fracture risk in clinical settings. For PA, we assigned individuals three levels (low, medium, and high) score according to the International Physical Activity Questionnaire guidelines. SB was quantified as an integer value representing the combined hours spent driving, using a computer, and watching television.
Table 1: Comparison of different methods in identifying risk factors for bone mineral density. Estimate and $\widehat{\text{SE}}$ represent the estimated linear regression coefficient and its corresponding standard error for each covariates. $\widehat{\text{SE}}$ ratio indicates the ratio of the standard error for a specific method compared with the classical method. PA denotes physical activity, and SB refers to sedentary behavior. The bold font represent the method that gives the smallest $\widehat{\text{SE}}$ ratio for each covariate.
Biological sex Age PA SB Smoking Alcohol
Classical Estimate -0.616 -0.190 0.019 0.040 0.006 -0.008
$\widehat{\text{SE}}$ 4.20E-03 4.12E-03 4.07E-03 4.18E-03 4.13E-03 4.20E-03
$\widehat{\text{SE}}$ ratio 1 1 1 1 1 1
PPI Estimate -0.604 -0.201 0.009 0.043 -0.026 -0.002
$\widehat{\text{SE}}$ 5.73E-03 6.26E-03 6.35E-03 6.48E-03 6.39E-03 6.47E-03
$\widehat{\text{SE}}$ ratio 1.363 1.517 1.560 1.553 1.549 1.541
EIF*-based Estimate -0.605 -0.200 0.010 0.042 -0.023 -0.003
$\widehat{\text{SE}}$ 5.28E-03 5.74E-03 5.83E-03 5.95E-03 5.87E-03 5.93E-03
$\widehat{\text{SE}}$ ratio 1.257 1.391 1.433 1.424 1.423 1.412
PSPA Estimate -0.614 -0.181 0.010 0.040 -0.008 -5.15E-07
$\widehat{\text{SE}}$ 3.74E-03 3.74E-03 3.74E-03 3.82E-03 3.80E-03 3.81E-03
$\widehat{\text{SE}}$ ratio 0.892 0.907 0.919 0.914 0.920 0.908
We regressed DXA-BMD on these variables using data from the UK Biobank (UKB). In the UKB, DXA-BMD measurements are available for only 10% of the participants. Therefore, we employed the Softimpute algorithm to impute DXA-BMD values for the remaining 90% individuals in the unlabeled dataset. To prevent overfitting, cross-validation was applied for the imputation in the labeled data. We consider classical, PPI, and EIF*-based method as baseline method.
Table 1 presents the inference results. Both the PPI and EIF*-based methods exhibit larger standard errors across all covariates compared to the classical method. In contrast, PSPA shows smaller standard errors for all coordinates when compared to the classical method, demonstrating its feature for element-wise variance reduction.
In conclusion, the real data results are consistent with our theoretical and simulation results, demonstrating that our method PSPA provides more efficient inference results in real-world post-prediction inference applications.
8 Comparison with PPI++
Finally, we present a comparison between PPI++  Angelopoulos et al., 2023b and our method PSPA, both theoretically and empirically. First, we note that PPI++ can be reviewed as a special case of PSPA when each element of the weighting vector ${\bm{\omega}}$ is constrained to the same values. PSPA  degenerates to PPI++  for one-dimensional estimation task such as mean estimation. Moreover, PPI++ cannot be applied to the case where the covariates are predicted by ML instead of the labels. Theoretically, since PPI++ can be reviewed as a special case of PSPA, PPI++ is less efficient than PSPA by Proposition 1. Moreover, PPI++ fails to guarantee element-wise variance reduction since it only incorporates one scalar for variance reduction.
Empirically, we compared PPI++ with PSPA for all experiments detailed in this paper. We present the simulation results in Figure 5 and real data applications in Figure 6 and Table 2, respectively. Simulation results show that both PPI++ and PSPA have the correct confidence interval coverage. However, PSPA has narrower confidence intervals compared with PPI++. In real data applications, PSPA identifies more sex-biased gene expressions than PPI++ in the GTEx example and has a smaller estimated standard error compared to PPI++  for every covariates in the DXA-BMD example. These results are consistent with Proposition 1, showing PSPA is more efficient than PPI++ in post-prediction inference applications.
@@ 9 Discussion
We have provided a simple yet powerful method, PSPA, to improve the efficiency of statistical inference with arbitrary ML predictions. To enable inference, we establish the consistency and asymptotic normality of the proposed estimator and prove its superiority over baseline methods. Through extensive simulations and real data applications, we demonstrate the superiority of the PSPA estimator over alternatives.
This work fits into a variety of scientific problems and unlocks several directions for future research. One can apply the same principle to more complicated data structures and missing mechanisms, for example in cases where a multi-dimensional outcome or covariate is subject to missing. It would be an interesting future work to carefully discuss relationships between ${\bf X}$ and ${\bf Z}$ and how they may affect the magnitude of the efficiency improvement.
The R codes to implement PSPA, benchmark methods, and replicate the simulation and real data analysis, is available at https://github.com/qlu-lab/pspa.
@@ Acknowledgments
Appendix A Supplementary Figures and Tables
Figure 5: Comparison of PPI++ and PSPA in simulation. This figure shows the coverage of the confidence interval and the relative ratio of the width of the confidence interval compared to the classical method for linear and logistic regression. ML is used to predict the labels. Panels (a)-(d) show the coverage of the confidence interval. Panels (e)-(h) show the relative ratio of the width of the confidence interval in comparison with the classical method. Panels (a), (b), (e), and (f) correspond to settings with varying sample sizes of unlabeled data. Panels (c), (d), (g), and (h) correspond to settings with different levels of imputation accuracy. The dashed line represent $y=0.95$ in (a)-(d) and $y=1$ in (e)-(h).
Figure 6: Comparison of PPI++ and PSPA in identifying sex-biased genes in GTEx.
Table 2: Comparison of PPI++ and PSPA in identifying risk factors for bone mineral density. Estimate and $\widehat{\text{SE}}$ represent the estimated linear regression coefficient and its corresponding standard error for each covariate. $\widehat{\text{SE}}$ ratio indicates the ratio of the standard error for a specific method compared with the classical method. PA denotes physical activity, and SB refers to sedentary behavior. The bold font represents the method that gives the smallest $\widehat{\text{SE}}$ ratio for each covariate.
Biological sex Age PA SB Smoking Alcohol
PPI++ Estimate -0.613 -0.193 0.016 0.041 -0.002 -0.006
$\widehat{\text{SE}}$ 3.86E-03 3.79E-03 3.84E-03 3.91E-03 3.87E-03 3.89E-03
$\widehat{\text{SE}}$ ratio 0.920 0.919 0.943 0.935 0.939 0.926
PSPA Estimate -0.614 -0.181 0.010 0.040 -0.008 -5.15E-07
$\widehat{\text{SE}}$ 3.74E-03 3.74E-03 3.74E-03 3.82E-03 3.80E-03 3.81E-03
$\widehat{\text{SE}}$ ratio 0.892 0.907 0.919 0.914 0.920 0.908